Ann and Nancy use a metal alloy that is 17.3% copper to make jewelry. how many ounces of a 17% alloy must be mixed with a 18% alloy to form 80 ounces of the desired alloy?
Answer by mananth(16946)

percent ----------------quantity
Alloy I 17 ----------------x ounces
Alloy II 18 ----------------80-x ounces
Total 17.30% ---------------- 80 ounces
17x+18(80-x)=17.3*80
17x+1440-18x=1384
17x-18x =1384-1440
-x=-56
x=56 ounces 17.00%
24 ounces 18.00%